OpenStudy (anonymous):

just need to see if my AP calc answer is right will post question. medal for best response.

OpenStudy (anonymous):

\[\int\limits_{0}^{x^{2}} \sqrt{1+r^{3}}\] use FTC to integrate

OpenStudy (anonymous):

i got\[\frac{ 2 }{ 3}\sqrt{1+x ^{2}}\] can anyone second that?

OpenStudy (mathmale):

I don't have a reference book with me to double check, but still see enough in the problem you've posted to ask whether this problem asks you to DIFFERENTIATE (not integrate) a function which is defined as the (unfinished) integral of some function.

OpenStudy (anonymous):

lol wut

OpenStudy (anonymous):

looks like it says to use the fundamental theorem of calculus F(b)-F(a)

OpenStudy (anonymous):

I'm just wondering if what i got for an answer is right or wrong

OpenStudy (mathmale):

If you want to understand the math involved, I'm your man. If you just want the answer to a problem, or want someone to give you a "right" or "wrong" response, I'm not.

OpenStudy (anonymous):

alright lol thanks anyway man. you probably know more than me ahah.

OpenStudy (anonymous):

that is what i got, too

OpenStudy (anonymous):

lol cool finally